OpenStudy (anonymous):

Can someone check this over to see if my answer is accurate? 12(x+2)-10-4x=40 I had to find the value of x and for the value I got was 6

OpenStudy (anonymous):

Sure one sec

OpenStudy (anonymous):

Kk @Etrainx

OpenStudy (anonymous):

Hm, I got something different, let me go over it one more time.

OpenStudy (anonymous):

Kk @Etrainx

OpenStudy (anonymous):

Well here is what I did, 12(x+2) becomes 12x + 24. so that means that 12x+24-10-4x=40 12x-4=8x 24-10 = 14 40-14=26 26/8= 3.25

OpenStudy (anonymous):

Yeah I made error while solving it while distributing the 12 to the 2 i didnt multiply it to get 12x+24

OpenStudy (anonymous):

Ah ok, well there you go :D

OpenStudy (anonymous):

Are u a mathematician ?

OpenStudy (anonymous):

Nah, just signed up today, needed some help with my summer work trig packet.

OpenStudy (anonymous):

@Etrainx I signed up yesterday I need so much help with my math summer homework